Find the probabilty if all the letters of the word "EQUATION" are arranged that first letter is a viowel and last is a consonant.

Grade:11

1 Answers

Neer Varshney
13 years ago

5 vowels and 3 consonants in this word.

Therefore number of ways= 5*6*5*4*3*2*1*3 = 6!*15.

Total word possibles are 8!.

Probability = 6!*15 / 8! = 15/56.